OpenStudy (anonymous):

Calculus f(x) = ( 20)/(x^2 + 21 ) (1)Find f'(x) = (2)Find f''(x) = (3)Find the open interval(s) where the function is increasing: (4)Find the open interval(s) where the function is decreasing: (5)Find the open interval(s) where the function is concave up: (6)Find the open interval(s) where the function is concave down:

ganeshie8 (ganeshie8):

what have u tried so far

OpenStudy (anonymous):

I found F' and F'' After that I lose myself with the open intervals, I missed out on class those two days. I've been looking in the calculus book and I have no clue where to start, or which principle to use here.

ganeshie8 (ganeshie8):

thats a good start, so what do u have for f'(x) and f''(x) ?

ganeshie8 (ganeshie8):

use this for parts 3 and 4 : f'(x) > 0 : function is increasing f'(x) < 0 : function is decreasing

ganeshie8 (ganeshie8):

you just need first derivative for those parts

OpenStudy (anonymous):

f'(x)= -(40 x)/(x^2+21)^2 f''(x)=(120 (x^2-7))/(x^2+21)^3 Do I use f' or f''?

ganeshie8 (ganeshie8):

f'

OpenStudy (anonymous):

just plug a zero into f'(x) then? or a -1 and 1?

ganeshie8 (ganeshie8):

\(\large f'(x) = \dfrac{-40x}{(x^2+21)^2}\) notice that the denominator is always positive since it is a square of something

ganeshie8 (ganeshie8):

the numerator is negative if x is positive and is positive if x is negative right ?

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

that means f'(x) > 0 : when x is less than 0 and f'(x) < 0 : when x is greater than 0

ganeshie8 (ganeshie8):

so can we say : the function is increasing in interval (-infty, 0) the function is decreasing in (0, infty) ?

OpenStudy (anonymous):

Yes, we can, now my question is how do we know that the function is increasing in interval (-infty, 0) and the function is decreasing in (0, infty) from what is given? Does this come from a specific principle?

ganeshie8 (ganeshie8):

recall that the first derivative at a point represents the slope of a tangent line at that point

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

notice that the graph is increasing for the left half part and decreasing afterwards

ganeshie8 (ganeshie8):

lets draw few tangents

OpenStudy (anonymous):

Sounds good to me

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

clearly the tangents are rising when the graph is increasing, so : \[\large \text{slope of tangents for left part is positive}\]

OpenStudy (anonymous):

for the right its negative

ganeshie8 (ganeshie8):

you know that slope of tangent at a point is same as the first derivative

ganeshie8 (ganeshie8):

yep you get the idea

OpenStudy (anonymous):

Ok Got it! which means that the answer to question 5 would be DNE correct?